# 2cm2cm1cm1cm

### v1

@@ 1 Introduction
Off-diagonal interaction terms of the form
@@
In the present paper, for a fairly general function $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ we study the effect of the off-diagonal interaction term
@@
where $b_{x}$ is the annihilation operator of a spinless boson on the site $x$ of the two-dimensional (square) lattice $\mathbb{Z}^{2}$, while $c_{y}$ represents the annihilation of a fermion pair of zero total spin, the two components of which are spread around the lattice position $y\in\mathbb{Z}^{2}$. The purely fermionic part of the considered model corresponds to the extended Hubbard Hamiltonian, as used in the context of ultracold atoms, ions, and molecules [11], while the purely bosonic component refers to an ideal gas, i.e., it has only a kinetic part (or “hopping term”), without interbosonic interactions. Because of the fermionic part, which is not exactly diagonalizable, the behavior of the full quantum many-body system, outside perturbative regimes, is almost inaccessible with the mathematical tools at our disposal. We thus consider only a three-body problem, by restricting the model to the sector of one boson and two fermions of opposite spins. In fact, the system restricted to this particular sector is very interesting, both mathematically and physically.
@@
Physically, the model is related to cuprate superconductors, like for instance $\mathrm{La}_{2-x}\mathrm{Sr}_{x}\mathrm{CuO}_{4}$ (LaSr 214) and $\mathrm{La}_{2-x}\mathrm{Ba}_{x}\mathrm{CuO}_{4}$. It is known [12, 13, 14] that in such crystals charge transport occurs within two-dimensional isotropic layers of copper oxides. This is why we consider here quantum particles on $2$-dimensional lattices $\mathbb{Z}^{2}$. A convincing microscopic mechanism behind superconductivity at high critical temperature is still lacking even after almost four decades of intensive theoretical and experimental studies. Many physicists believe that the celebrated Hubbard model could be pivotal, one way or another, in order to get a microscopic theory of high-temperature superconductivity, but many alternative explanations or research directions have also been considered in theoretical physics. For some of the more popular models for cuprate superconductors, see, e.g., [13, Chap. 7].
@@
In many theoretical approaches to this problem, the existence of polaronic quasiparticles in relation with the very strong Jahn-Teller (JT) effect associated with copper ions is neglected, as stressed in [15, Part VII]. The role of polarons is however highlighted in [16], since the JT effect actually led to the discovery of superconductivity in cuprates in 1986. See [17, p. 2] or [18, 19]. Our theoretical approach differs from most popular ones, being based on the existence of JT bipolarons in copper oxides, as is discussed in the literature [20] at least as early as 1990. The physics behind this approach is explained in detail in [21], where a simplified version of the model studied here is considered. In our microscopic model for cuprate superconductors, as presented in [22, 21], the bosonic operator $b_{x}$ ($b_{x}^{\ast}$) in (4) refers to the annihilation (creation) of a JT bipolaron, whereas the fermionic one $c_{y}$ ($c_{y}^{\ast}$) annihilates (creates) a fermion pair, which is reminiscent of Cooper pairs in conventional superconductivity.
@@
As in the present paper, no ad hoc assumptions, in particular concerning anisotropy, are made in [21]. In fact, [21] proves that unconventional pairing may occur, breaking spontaneously discrete symmetries of the model, like the $d$-wave pairing, whose wave function is antisymmetric with respect to $90^{\circ}$-rotations. It turns out that electrostatic (screened Coulomb) repulsion is crucial for such unconventional pairings, which are meanwhile shown to be concomitant with a strong depletion of superconducting pairs. Notice that the results of [21] are coherent with experimental observations on the cuprate LaSr 214: The coherence length at optimal doping and the $d$-wave pair formation in the pseudogap regime, i.e., at temperatures much higher than the superconducting transition temperature, are predicted in good accordance with experimental data.
@@
This was already done in [21, 22], but only for usual (non-extended) Hubbard interactions and one-range creation / annihilation operators of fermion pairs. Even in this specific case, the dispersion relation of dressed bound fermion pairs was analyzed only to a level of detail enough to deduce unconventional pairings near the ground state. By contrast, in the present paper other important properties of the dispersion relation, like its regularity, are studied for the first time and in a more general framework. Using Combes-Thomas estimates we show, among other things, that the dressed bound fermion pairs are localized, in the sense that the fermion-fermion correlation decays very fast in space. Group velocities and tensor masses of dressed bound fermion pairs are also shown to exist under very natural conditions on the (absolutely summable function) $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ appearing in (4). In fact, our analysis allows one to accurately understand which features of the exchange strength function $\upsilon$ can strengthen the stability of the dressed bound fermion pairs. For instance, $\upsilon$ has to be sufficiently strong and localized in Fourier space in order to get a sufficiently strong “gluing effect”. Additionally, the boson should be heavier than two fermions. Notice that this second condition is consistent with the physical interpretation that the boson is a bipolaron, which is known to be (effectively) much heavier than the fermions (electrons or holes), in superconducting cuprates. Observe additionally that the very large mass of bipolarons (and polarons, in general), is one of the main arguments used to discredit theoretical approaches based on bipolarons, because it is known from experiments that the charge carriers in superconducting cuprates have an effective mass comparable to that of electrons and holes. In fact, we prove that the effective mass of bound pairs mainly depends on the properties of the function $\upsilon$, that encodes the fermion-boson exchange processes, but not much on the mass of the boson itself. This issue is discussed in [21], in detail. See also the discussion at the end of Section 3.2. That is why we are interested in results concerning the mass tensor for bound pairs and we think we provide here a convincing solution for the “mass paradox” related to bipolaronic pairing mechanisms in the microscopic theory of cuprate superconductors.
@@
We also study here scattering properties of the three-body model in two channels: the unbound and bound pair channels. The first one corresponds to the case in which two fermions start far apart from each other and only experience a very weak repulsion force due to the extended Hubbard interaction, while the probability that they bind together to form a boson is very small. In this situation, we show that two (almost) freely propagating fermions in the distant past can come together and interact with one another, either via the repulsive electrostatic force or by exchanging a boson, and then propagate away, again freely in the distant future. In this channel, the scattering matrix can be explicitly computed via convergent (Dyson) series, making in particular the study of the scattering effect of the fermion-boson-exchange interaction (1) technically uncomplicated.
@@
The bound pair (scattering) channel corresponds to the case in which dressed bound fermion pairs are formed. In contrast with the first channel, now there is a non-negligible bosonic component related with the exchanged boson that “glues” the two fermions together. We prove that those (spatially localized) dressed bound fermion pairs effectively move like a free (quantum spinless) particle. In this case, strictly speaking in the physical sense, there is no scattering and the pairs evolve freely in space, governed by an effective dispersion relation, the Fourier transform of which is the effective hopping strength for the (spatially localized) dressed bound pairs.
@@
We additionally prove that all these properties hold also true in the limit of large on-site fermionic repulsions, provided that two fermions on two different lattice sites can interact via the fermion-boson exchange interaction. It refers to a hard core limit, preventing two fermions from occupying the same lattice site. For cuprate superconductors, it is an important issue addressed and answered here, because of the undeniable experimental evidence of very strong on-site Coulomb repulsions in cuprates, leading to the universally observed Mott transition at zero doping [23, 24].
@@
To conclude, the mathematical properties of the model studied in the present work are well understood and, as a consequence, the model can serve as a prototypical example of a quantum system including exchange interaction terms of the form (1). From a physics viewpoint, it is also interesting, since dressed bound fermion pairs are good candidates for superconducting charge carriers in cuprate superconductors, as advocated in [21]. Our main results are Theorems 3.1, 3.5, 3.6, 3.9, 3.11 and 3.14. The paper is organized as follows: Section 2 explains in detail the model, while Section 3 gives the main results. Technical outcomes, along with all their proofs, are gathered in Section 4. Section 5 is an appendix collecting some important, albeit well-known, mathematical results used here, making the paper self-contained.

### v2

@@ 1 Introduction
1.1 Exchange Interactions and High-Tc Superconductivity
Exchange interactions in Mathematical Physics. Off-diagonal interaction terms of the form
@@ 1.1 Exchange Interactions and High-Tc Superconductivity
Three-body fermion-boson exchange interactions. In the present paper, for a fairly general function $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ we study the effect of the off-diagonal interaction term
@@
where $b_{x}$ is the annihilation operator of a spinless boson on the site $x$ of the two-dimensional (square) lattice $\mathbb{Z}^{2}$, while $c_{y}$ represents the annihilation of a fermion pair of zero total spin, the two components of which are spread around the lattice position $y\in\mathbb{Z}^{2}$. See Fig. 1.
[…]
Figure 1: Illustration of fermion-boson exchange interactions in the form of two Feynman diagrams. In theoretical physics, a Feynman diagram visually represents the mathematical expressions that describe the behavior and interactions of quantum particles. In the example on the left, the two arrows indicate that two fermions, named $(f)$, “collide” to create a new particle, the boson $(b)$. The oscillating line is generally used to describe an interaction with a mediator, which can be seen by combining the two diagrams: two fermions $(f)$ interact to produce a boson, which annihilates again to produce two fermions $(f)$. This can lead to an effective interaction between fermions. In particular, this process could produce a pair of fermions ($f-f$) bonded by the exchange of a bosonic field ($b$), according to the coupling function $\upsilon$. This is typically what we are going to show. Note that the opposite combination can also be made: a boson $(b)$ is destroyed to create two fermions $(f)$, which annihilate to recreate a boson $(b)$. This does not really create an interaction as such, but a kinetic term, or seen another way, a self-interaction on the boson $(b)$. The combination of two diagrams refers to a perturbative approach of second order, but we can also combine several of the same diagrams (perturbative approach of order $n$). Note, however, that no such perturbative argument is used here.
Note that the opposite combination can also be made: a boson $b$ is destroyed to create two fermions $f$, which annihilate to recreate a boson $b$. This does not really create an interaction as such, but a kinetic term, or seen another way, a self-interaction. The combination of two diagrams refers to a perturbative approach of second order, but we can also combine several of the same diagrams (perturbative approach of order $n$).
@@
The purely fermionic part of the considered model corresponds to the extended Hubbard Hamiltonian, as used in the context of ultracold atoms, ions, and molecules [11], while the purely bosonic component refers to an ideal gas, i.e., it has only a kinetic part (or “hopping term”), without interbosonic interactions. Because of the fermionic part, which is not exactly diagonalizable, the behavior of the full quantum many-body system, outside perturbative regimes, is almost inaccessible with the mathematical tools at our disposal.
We thus consider only a three-body problem, by restricting the model to the sector of one boson and two fermions of opposite spins. In fact, the system restricted to this particular sector is very interesting, both mathematically and physically. Note that such sector restrictions in Fock spaces are also performed for the study of the Pauli-Fierz and Nelson models [26, 28, 27] in non-relativistic Quantum Field Theory (QFT).
Physical context: High-$T_{c}$ superconductivity of cuprates. Physically, the model is related to cuprate superconductors [^2] You can take for instance the cuprate $\mathrm{La}_{2}\mathrm{CuO}_{4}$, which is a Mott insulator with an antiferromagnetic phase at low temperature. As with semiconductors, it is doped with atoms like $\mathrm{Sr}$ or $\mathrm{Ba}$, which add a few charge carriers (in this case, holes). Then, with moderate doping $x$, the material becomes superconducting at low temperatures. This is the meaning of the chemical formulae $\mathrm{La}_{2-x}\mathrm{Sr}_{x}\mathrm{CuO}_{4}$ and $\mathrm{La}_{2-x}\mathrm{Ba}_{x}\mathrm{CuO}_{4}$, $x$ being a small number characterizing the cuprate doping. See also Section 5.1., like for instance $\mathrm{La}_{2-x}\mathrm{Sr}_{x}\mathrm{CuO}_{4}$ (LaSr 214 or LSCO) and $\mathrm{La}_{2-x}\mathrm{Ba}_{x}\mathrm{CuO}_{4}$. It is known [12, 13, 14] that in such crystals charge transport occurs within two-dimensional isotropic layers of copper oxides. This is why we consider here quantum particles on $2$-dimensional lattices $\mathbb{Z}^{2}$.
@@
A convincing microscopic mechanism behind superconductivity at high critical temperature is still lacking even after almost four decades of intensive theoretical and experimental studies. See Section 5.1 for more details. Many physicists believe that the celebrated Hubbard model could be pivotal, one way or another, in order to get a microscopic theory of high-temperature superconductivity, but many alternative explanations or research directions have also been considered in theoretical physics. For some of the more popular models for cuprate superconductors, see, e.g., [13, Chap. 7].
In many theoretical approaches to this problem, the existence of polaronic quasiparticles in relation with the very strong Jahn-Teller (JT) effect associated with copper ions is neglected, as stressed in [15, Part VII]. The role of polarons is however highlighted in [16], since the JT effect actually led to the discovery of superconductivity in cuprates in 1986. See [17, p. 2] or [18, 19].
@@
Our theoretical approach differs from most popular ones, being based on the existence of JT bipolarons in copper oxides, as is discussed in the literature [20] at least as early as 1990. The physics behind this approach is explained in detail in [21], where a simplified version of the model studied here is considered. In our microscopic model for cuprate superconductors, as presented in [22, 21], the bosonic operator $b_{x}$ ($b_{x}^{\ast}$) in (4) refers to the annihilation (creation) of a JT bipolaron, whereas the fermionic one $c_{y}$ ($c_{y}^{\ast}$) annihilates (creates) a fermion pair, which is reminiscent of Cooper pairs in conventional superconductivity.
@@
Bipolaronic pairing mechanisms and cuprate superconductivity. As in the present paper, no ad hoc assumptions, in particular concerning anisotropy, are made in [21]. In fact, [21] proves that unconventional pairing may occur, breaking spontaneously discrete symmetries of the model, like the $d$-wave pairing, whose wave function is antisymmetric with respect to $90^{\circ}$-rotations. It turns out that electrostatic (screened Coulomb) repulsion is crucial for such unconventional pairings, which are meanwhile shown to be concomitant with a strong depletion of superconducting pairs.
Notice that the results of [21] are coherent with experimental observations on the cuprate LaSr 214: The coherence length at optimal doping and the $d$-wave pair formation in the pseudogap regime, i.e., at temperatures much higher than the superconducting transition temperature, are predicted in good accordance with experimental data. In addition to the $d$-wave pairing and the high-temperature pseudogap regime, the model considered here also captures another very special feature of high-Tc cuprate superconductors, namely the density waves [23]. For more details, see also Section 5.1.
In fact, it is shown in [22, Section 4.1] that three-body fermion-boson exchange interactions, like the one studied in this paper, imply an effective fermion-fermion interaction. Then, by considering the mean-field limit of it, which corresponds to taking couplings (4) that are very localized in momentum space [22, Section 4.2], it was rigorously proven [24] that, below the critical temperature, the equilibrium states of the (purely fermionic) associated many-body Hamiltonian exhibit periodic modulation in space of the charge density, even incommensurate with respect to the lattice spacing.
1.2 Mathematical Results
Previous results. To our knowledge, the model considered here has not been studied mathematically, apart from our own articles [21, 22] published in recent years. See also the Ph.D. thesis [24]. Mathematical studies for explicit exchange interaction terms are mainly those presented above. As far as we know, concerning its physical interpretation regarding cuprate superconductivity, our approach has also never been considered by other physicists and we therefore doubt that any theoretical results in this direction exist in the literature. For more details, see the introductory discussions in [21], which give a concise overview of theories of high-temperature superconductivity.
@@ 1.2 Mathematical Results
This was already done in [21, 22], but only for usual (non-extended) Hubbard interactions and one-range creation / annihilation operators of fermion pairs. Even in this specific case, the dispersion relation of dressed bound fermion pairs was analyzed only to a level of detail enough to deduce unconventional pairings near the ground state. By contrast, in the present paper other important properties of the dispersion relation, like its regularity, are studied for the first time and in a more general framework.
Last but not least, the localization of dressed bound fermion pairs or the scattering properties of the model have not been studied before.
Localized dressed bound fermion pairs. Using Combes-Thomas estimates we show, among other things, that the dressed bound fermion pairs are localized, in the sense that the fermion-fermion correlation decays very fast in space. Group velocities and tensor masses of dressed bound fermion pairs are also shown to exist under very natural conditions on the (absolutely summable function) $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ appearing in (4).
In fact, our analysis allows one to accurately understand which features of the exchange strength function $\upsilon$ can strengthen the stability of the dressed bound fermion pairs. For instance, $\upsilon$ has to be sufficiently strong and localized in Fourier space in order to get a sufficiently strong “gluing effect”. Additionally, the boson should be heavier than two fermions.
Notice that this second condition is consistent with the physical interpretation that the boson is a bipolaron, which is known to be (effectively) much heavier than the fermions (electrons or holes), in superconducting cuprates. Observe additionally that the very large mass of bipolarons (and polarons, in general), is one of the main arguments used to discredit theoretical approaches based on bipolarons, because it is known from experiments that the charge carriers in superconducting cuprates have an effective mass comparable to that of electrons and holes.
In fact, we prove that the effective mass of bound pairs mainly depends on the properties of the function $\upsilon$, that encodes the fermion-boson exchange processes, but not much on the mass of the boson itself. This issue is discussed in [21], in detail. See also the discussion at the end of Section 3.2. That is why we are interested in results concerning the mass tensor for bound pairs and we think we provide here a convincing solution for the “mass paradox” related to bipolaronic pairing mechanisms in the microscopic theory of cuprate superconductors.
Relationship with the enhanced binding of QFT. The formation of dressed bonded fermion pairs as described above is reminiscent of what is known as enhanced binding in Quantum Field Theory (QFT). For more details on this phenomenon we recommend the lecture notes [25], where it is well explained in the context of non-relativistic QFT. See also the references therein.
For example, the Pauli-Fierz model, which refers to non-relativistic quantum charge particles interacting with a massless quantized radiation field (photons), can have at low energies a dressed particle with an effective mass bigger than the non-interacting one, leading to the existence of a ground state for the model. A similar fact occurs in the Nelson model, in which $N$ quantum particles interact linearly with a field of photons (or mesons). The formation of such dressed particles is a direct consequence of the bosonic field acting as mediator of a force.
Indeed, in both cases, the model involves a sum of interaction terms of the form $\psi_{k}\otimes b_{k}+\bar{\psi}_{k}\otimes b_{k}^{\ast}$, coupling the $N$-body quantum system with a spinless boson field of momentum $k$ via annihilation/creation operators $b_{k},b_{k}^{\ast}$. Note that in this case there is no transformation of particles of one type into another, as in the exchange interactions described above, but both cases are still similar, especially as we are carrying out our analysis in the sector with only two fermions and one boson. This makes the comparison quite relevant, even if the model and mathematical methods considered here have essential differences as compared to the previous ones.
Scattering properties of the model. We also study here scattering properties of the three-body model in two channels, the unbound and bound pair channels:
  • The unbound pair (scattering) channel corresponds to the wave and scattering operators with respect to fermionic part, respectively defined via the strong limits
    $W^{\pm}\doteq s-{\lim\limits_{t\rightarrow\pm\infty}}\mathrm{e}^{it\mathrm{H}}\mathrm{e}^{-it\mathrm{H}_{f}}P_{\mathrm{ac}}\left(\mathrm{H}_{f}\right)\text{\qquad and\qquad}S\doteq\left(W^{+}\right)^{\ast}W^{-}\ ,$
@@
    where $\mathrm{H}_{f}$ is a generic, purely fermionic Hamiltonian representing free fermions that do not interact with any bosonic field, $\mathrm{H}$ is the Hamiltonian of the full model and $P_{\mathrm{ac}}(\mathrm{H}_{f})$ is the orthogonal projection onto the absolutely continuous space of $\mathrm{H}_{f}$. It refers to the case in which two fermions start far apart from each other and only experience a very weak repulsion force due to the extended Hubbard interaction, while the probability that they bind together to form a boson is very small. It In this situation, we show that two (almost) freely propagating fermions in the distant past can come together and interact with one another, either via the repulsive electrostatic force or by exchanging a boson, and then propagate away, again freely in the distant future. In this channel, the scattering matrix can be explicitly computed via convergent (Dyson) series, making in particular the study of the scattering effect of the fermion-boson-exchange interaction (1) technically uncomplicated.
  • The bound pair (scattering) channel corresponds to the time evolution $\mathrm{e}^{it\mathrm{H}}\mathfrak{P}$, $t\in\mathbb{R}$, where $\mathrm{H}$ is again the Hamiltonian of the full model and $\mathfrak{P}$ is an isometry from the $L^{2}$-functions on the Brillouin zone to the subspace associated with the fiber bound states of $\mathrm{H}$. We show in particular that
    $\mathrm{e}^{it\mathrm{H}}\mathfrak{P}=\mathfrak{P}\mathrm{e}^{itM_{\mathrm{E}\left(\cdot\right)}}\ ,\qquad t\in\mathbb{R}\ ,$
    with $M_{\mathrm{E}\left(\cdot\right)}$ being some multiplication operator given by the dispersion relation $k\mapsto\mathrm{E}\left(k\right)$ characterizing the (fiber) bound states at fixed quasi-momentum in the normalized Brillouin zone $\mathbb{T}^{2}\doteq[-\pi,\pi)^{2}$. In terms of wave operators, it follows that
    $W^{\pm}\doteq s-{\lim\limits_{t\rightarrow\pm\infty}}\mathrm{e}^{it\mathrm{H}}\mathfrak{P}\mathrm{e}^{-itM_{\mathrm{E}\left(\cdot\right)}}P_{\mathrm{ac}}\left(M_{\mathrm{E}\left(\cdot\right)}\right)=\mathfrak{P}P_{\mathrm{a}\mathrm{c}}\left(M_{\mathrm{E}\left(\cdot\right)}\right)\ ,$
    which gives a scattering operator equal to
    $S\doteq\left(W^{+}\right)^{\ast}W^{-}=P_{\mathrm{a}\mathrm{c}}\left(M_{\mathrm{E}\left(\cdot\right)}\right)\ .$
@@
    It refers to the case in which dressed bound fermion pairs are formed. In contrast with the first channel, now there is a non-negligible bosonic component related with the exchanged boson that “glues” the two fermions together. We prove that those (spatially localized) dressed bound fermion pairs effectively move like a free (quantum spinless) particle. In this case, strictly speaking in the physical sense, there is no scattering and the pairs evolve freely in space, governed by an effective dispersion relation, the Fourier transform of which is the effective hopping strength for the (spatially localized) dressed bound pairs.
@@
Composite system at strong on-site Hubbard repulsions. We additionally prove that all these properties hold also true in the limit of large on-site fermionic repulsions, provided that two fermions on two different lattice sites can interact via the fermion-boson exchange interaction. It refers to a hard core limit, preventing two fermions from occupying the same lattice site.
For cuprate superconductors, it is an important issue addressed and answered here, because of the undeniable experimental evidence of very strong on-site Coulomb repulsions in cuprates, leading to the universally observed Mott transition at zero doping [29, 30].
1.3 Concluding Remarks and Structure of the Paper
@@
To conclude, the mathematical properties of the model studied in the present work are well understood and, as a consequence, the model can serve as a prototypical example of a quantum system including exchange interaction terms of the form (1). From a physics viewpoint, it is also interesting, since dressed bound fermion pairs are good candidates for superconducting charge carriers in cuprate superconductors, as advocated in [21].
More specifically, our main results are Theorems 3.1, 3.5, 3.6, 3.9, 3.11 and 3.14. The paper is organized as follows: Section 2 explains in detail the model, while Section 3 gives the main results. Technical outcomes, along with all their proofs, are gathered in Section 4. Section 5 is an appendix that gathers important standard mathematical results used here, an overview of cuprate physics for non-physicists, as well as the Fock-space formalism, in order to make the article self-contained and accessible to a wide audience.
Remark 1.1 ($d$-dimensional lattices)
Our study focuses on two-dimensional lattice systems because of their application to the superconductivity of cuprates and, in particular, their $d$-wave symmetry. However, it can also be done at arbitrary dimension $d\geq 1$ provided the coupling functions used, i.e., $\mathrm{u},\mathfrak{p}_{1},\mathfrak{p}_{2},\upsilon:\mathbb{Z}^{d}\rightarrow\mathbb{R}_{0}^{+}$ below, stays absolutely summable. It is also important that the Fourier transforms $\hat{\upsilon}$, $\mathfrak{\hat{p}}_{1}$ and $\mathfrak{\hat{p}}_{2}$ of the functions $\upsilon$, $\mathfrak{p}_{1}$ and $\mathfrak{p}_{2}$ remain real-valued [^3] A real-valued absolutely summble function $f$ on $\mathbb{Z}^{d}$ has a real-valued Fourier transform iff $f(-z)=f(z)$. Considering two-dimensional systems that are invariant under $90^{\circ}$-degree rotations (like we did, because of cuprates), this property is always true and has not to be additionally imposed. continuous functions on the $d$-dimensional torus $\mathbb{T}^{d}$.